OpenStudy (anonymous):

help! what is the vertex of the graph of y=5(x+4)^2

OpenStudy (anonymous):

do you know what an equation in vertex form is?

OpenStudy (anonymous):

no how do i change it?

OpenStudy (anonymous):

its already in vertex form meaning you can pull the vertex out of the equation.

OpenStudy (anonymous):

Vertex = (h,k) vertex form is y= a(x-h)^2+k

OpenStudy (anonymous):

i know how to get the vertex out if the equation was something like y=x2 +bx+c but i dont know how for this one

OpenStudy (anonymous):

oops i wrote the equation wrong its y=5(x+4)^2 +3

OpenStudy (anonymous):

ok so based off this: Vertex = (h,k) vertex form is y= a(x-h)^2+k what do you think the vertex is?

OpenStudy (anonymous):

(4, 3)

OpenStudy (anonymous):

close but no, remember to get (x+h)^2 from (x-h)^2 the h has to be negative.

OpenStudy (anonymous):

so (-4, 3)

OpenStudy (anonymous):

Exactly! Now you got it!